# Which Demographics do LLMs Default to During Annotation?

### v2

@@ Abstract
Johannes Schäfer, Aidan Combs, Christopher Bagdon, Jiahui Li, Nadine Probol, Lynn Greschner, Sean Papay, Yarik Menchaca Resendiz, Aswathy Velutharambath, Amelie Wührlthanks: These authors contributed equally., Sabine Weber,  and Roman Klinger Fundamentals of Natural Language Processing, University of Bamberg roman.klinger@uni-bamberg.de
@@ 1 Introduction
In some text annotation tasks, it is feasible to obtain an aggregated ground truth label, for instance in named entity annotation (Yadav and Bethard, 2018) or semantic role labeling (Shi et al., 2020). In other tasks, perhaps even in the majority of tasks, it is more obvious that annotator’s traits influence the label assignments, for instance in sentiment annotation (Liu, 2012), emotion annotation (Klinger, 2023; Plaza-del Arco et al., 2024), or personality profiling (Neuman, 2015). The diversity of annotations, conditioned on annotators’ profiles, has been recognized as an important variable to consider, instead of aggregating all labels to an adjudicated score, which might not correspond to any of the annotators (Plepi et al., 2022).
@@
With the advent of (instruction-tuned) large language models, automatic data annotation and zero or few shot predictions became more popular (Brown et al., 2020). However, language models do not provide the same diversity as human annotators do by simply prompting them in a zero-shot setup, which can lead to a lower performance of the labeling process (Bagdon et al., 2024). To mitigate this problem, it is important to understand whether large language models exhibit a default persona when acting as annotators and how this can be controlled to ensure greater diversity in annotations. One concern is that large language models may reflect biases present in their training data, which can disproportionately emphasize certain viewpoints. This raises the issue of potential marginalization of minority perspectives in the outputs generated by these models, making it crucial to address this bias for equitable representation in annotation tasks. One idea to address this limitation is socio-demographic prompting (Muscato et al., 2024), where we specify characteristics like age, gender, ethnicity or socio-economic status.
@@
Previous research has explored the effects of using prompts informed by demographic or cultural contexts on model predictions, but did not find consistent patterns (Beck et al., 2024; Mukherjee et al., 2024). We build on top of this work and particularly contribute in two directions of socio-demographic prompting in a large demographic-annotated data set, namely Popquorn (Pei and Jurgens, 2023) that has not undergone an analysis from the LLM-perspective yet. (1) We study if large language models default to a particular demographic, i.e., we evaluate if the LLM-based predictions more closely align with those of people of a particular demographic when not being conditioned. (2) Additionally, we utilize placebos as suggested by Mukherjee et al. (2024) – irrelevant information that should not affect the model’s output – to compare the impact of demographic prompts and evaluate the model’s prediction stability. More concretely, we answer the following questions.
@@
The remainder of the paper is structured as follows: We review related work in Section 2 and explain our data set choices and prompt setup in Section 3. Section 4 presents the results and Section 5 concludes and explains possible future work.
@@ 2 Related Work
This paper relates to various areas which we review in the following, namely biases in language models, perspectivism, and (socio-demographic) prompting.
@@ 2.1 Inherent Knowledge and Biases in LLMs
Large language models may be understood as databases that store information encoded in the training data (Petroni et al., 2019). These data is stored in a probabilistic manner, which allows these models to show sometimes unexpected generalization capabilities beyond the original instructions available in the training data, i.e., emergent abilities (Wei et al., 2022). It has also been shown that prompting can adapt parameters during inference, similar to how backpropagation works (von Oswald et al., 2023). Therefore, it is reasonable to view prompting as a form of programming, as it involves manipulating model behavior by crafting specific inputs to achieve desired outputs (Beurer-Kellner et al., 2023).
@@ 2.2 Perspectivism
Natural language processing for a long time focused on seemingly objective tasks such as parsing or named entity recognition. Therefore, it has been a standard approach to adjudicate annotations into a single gold standard (Stubbs, 2011). To do so, a set of methods for aggreggation that considers the distributions and disagreements have been developed (Paun et al., 2018). More recently, research moved towards acknowledging the importance of disagreements. Plank (2022) has been one of the first to discuss challenges and potential approaches to this problem. Recent work looked into methods to consider disagreements (Fleisig et al., 2024) and discusses ethical considerations (Valette, 2024). Nowadays, this field perspectivism developed dedicated shared tasks (Uma et al., 2021) and workshops (Abercrombie et al., 2024).
Disagreement is related to confidence of annotators (and models), an aspect that has received some attention. Baumler et al. (2023) made use of this property in the context of active learning, to understand which instances require annotations by multiple people in order to to understand the disagreements. Troiano et al. (2021); Baan et al. (2024) showed that annotator’s own confidence predicts inter-annotator agreement scores.
To understand disagreement in labeling better, more and more corpora are being published with more detailed information about the annotators. Troiano et al. (2023) annotated event reports for emotions and appraisals and collected demographic information, personality, and current emotional state of both the person who lived through the event and multiple annotators that read the event description. Plepi et al. (2022) studied the role of demographics, automatically extracted from the data, on the perception of social norms. Bizzoni et al. (2022) discuss the role of individual differences on the judgement of literary quality. Romberg (2022) integrates perspectivism in argument mining, by making explicit the subjective nature of argument interpretation. May et al. (2024) study the effect of demographics on the role of numbers in social judgements. Frenda et al. (2023) study how the perception of irony varies by nationality, employment status, student status, ethnicity, age, and gender. Sachdeva et al. (2022) measures aspects of hate speech, including sentiment, disrespect, insult, attacking/defending, humiliation, inferior/superior status, dehumanization, violence, genocide, and a 3-valued hate speech benchmark label. They study these variables under the condition of identity target groups and annotator demographics. Xu et al. (2023) look into disagreement of legal case outcome differences. Next to subjective tasks, perspectivism has also been considered in seemingly objective tasks, for instance in named entity recognition (Peng et al., 2024) and natural language inference (Gruber et al., 2024).
@@ 2.4 Socio-Demographic Prompting
Automatic annotation with language models is not a replacement for human annotation. Humans have previous world knowledge, experiences, and perspectives on a matter that differ individually. LLMs have difficulties replicating these differences, causing issues in annotation. For instance, Bagdon et al. (2024) showed that the diversity of annotations that is beneficial in comparative annotations by multiple people is not straight-forwardly replaced by multiple runs of a language model. Lee et al. (2023) focused on this aspect and found poor alignment of the distribution of labels predicted by LLMs with human annotations on natural language inference tasks.
@@ 3.1 Data Sets
We chose the Popquorn data for our experiments (Pei and Jurgens, 2023). In contrast to data used by other previous research we are aware of, these data has been particularly sampled for the study of annotator’s properties and their impact on annotation tasks. Therefore, these data render themselves as a straight-forward choice also for an LLM-based analysis. The original data stems from the Ruddit corpus (Hada et al., 2021) which has originally been annotated via best–worst scaling. BWS may lead to different label frequencies than rating scale annotations and is not appropriate for skewed distributions (Finn and Louviere, 1992; Louviere et al., 2015). The authors of Popquorn reannotated with 1–5 rating scales, which corresponds to our annotation setup. Further, it is noteworthy that the creators of the data set removed annotators with a particularly low annotation competency according to Mace (Hovy et al., 2013), to filter out potential random noise.
The Popquorn data set consists of 45,000 annotations from 1,484 annotators and includes information on annotators’ gender, race, age, occupation and education. It is partioned into multiple NLP tasks. We use data from the offensiveness and politeness rating tasks. To equalize the number of annotations per instance, we sample three annotations for each instance. When sampling, we omit the annotations where the annotator preferred not to disclose demographic information. We exclude data instances with fewer than three annotations remaining. The two subsets of the sample which we use in our experiments consist of 4,500 annotations of 1,500 instances for offensiveness ratings and 11,151 annotations of 3,717 instances for politeness ratings.
3.2 Prompt Setup
We use three different prompt setups, namely socio-demographically informed (SD) prompts, placebo-informed (P) prompts, and prompts without additional information (N prompts). We show the templates for these prompts in Table 1. The input to the LLM consist of a concatenation of the system message with the respective text to be classified. The values for the demographic attributes are taken from the demographic data of the annotators included in the sample. The sets of values for the placebo attributes are displayed in Table 2. During prompting, each placebo attribute value is randomly sampled from the respective set.
@@ 3.2 Prompt Setup
Table 2: Values of placebo attributes used in our P prompts.
@@ 3.3 Model Choice and Access
We use two LLMs, namely GPT-4o [^1] https://openai.com/index/hello-gpt-4o/ and Claude [^2] https://www.anthropic.com/news/claude-3-5-sonnet, claude-3-5-sonnet-20240620. We access both models via their respective APIs. The cost for using GPT-4o was \$54 and the cost for using Claude was \$60.
@@ 3.4 LLM Output Parsing
We transform the texts generated by the LLMs into a defined label set of five categories, corresponding to a rating scale of 1 to 5 for each of the two tasks. We used the Langchain StrOutputParser [^3] https://api.python.langchain.com/en/latest/output_parsers/langchain_core.output_parsers.string.StrOutputParser.html to interpret the output from GPT-4o as one of the designated labels, while the output from Claude was utilized directly.
There were few cases where the output could not be parsed. We encountered 22 such cases with GPT-4o and three with Claude. Of the failed cases with GPT-4o, 14 were due to the input text from the original data set being in Polish instead of English. This resulted in the model predicting labels in Polish, despite the prompt specifying an English-only label set. The remaining eight cases involved very short text instances or questions, which the model misinterpreted. In these cases, GPT-4o asked for further input instead of performing the classification task.
For Claude, the three instances of failure involved actor names in the text. This triggered the model’s copyright protection protocols, preventing it from reproducing or paraphrasing potentially copyrighted content. Given the total number of instances analyzed in our experiments, this failure rate is insubstantial. Consequently, we disregarded the output in these cases and assigned the labels “not offensive” or “not polite”.
@@ 3.5 Evaluation Settings
We conduct various analyses of the labels generated by the LLMs, utilizing the different prompt setups described in Section 3.2. First, we examine the extent to which LLM predictions align with the judgments of different human annotators by comparing the outputs generated from non-demographic (N) prompts to the multiple annotations available for each instance. Second, we examine the impact of socio-demographic prompting by comparing the models’ automatic annotations produced with SD prompts to those produced with N prompts. In this analysis, we assess the differences in the LLMs’ annotations when demographic data is included versus when it is omitted. Thirdly, we investigate the effect of placebo prompting in a similar manner by comparing the automatic annotations generated with P prompts to those created with N prompts.
@@ 4 Results
In this section, we present the results of our experiments based on the annotations generated by the LLMs for two labeling tasks. Each of our research questions is addressed with particular results derived from the various experiments conducted. Detailed results for individual prompts given in Section A.2 show that the two N prompts behave similarly. We also observed this for the other prompt types (see Table 1). Thus, we here report all results as an average over the two respective prompts for the different prompt types.
@@
We approach the identification of the default persona of models in two different ways following the first two settings as described in Section 3.5. The following sections present and discuss the results of these analyses.
@@
Which socio-demographic attributes of human annotators does an LLM inherently mimic in the absence of explicit information (N prompts)? Table 3 shows the results for this analysis which corresponds to the first setting described in Section 3.5. We combine socio-demographic attributes that are represented in only a few cases. The table specifies the reference categories (ref.) used for each of the categorical variables.
@@
Table 3 shows the coefficients of regression models predicting absolute values of the distance between NLLM annotations and annotations provided by human annotators for the same instance. Independent variables are the sociodemographic characteristics of the human annotators. We report results separately for the two models (GPT-4o and Claude) and two classification tasks. The reported coefficients can be interpreted as the effect of particular human demographic categories on the distance between human and LLM annotations (change in distance between the category in question and the reference category). Positive coefficients indicate that the respective LLM is further from the human annotators in that demographic category than human annotators in the reference category, i.e., less accurate. So, categories with positive, statistically significant coefficients are those categories which the respective model does not default to.
@@
This analysis (see Table 3) reveals significant biases in the LLMs’ predictions based on demographic factors, particularly concerning race and age. Specifically, the models demonstrate a tendency to align more closely with annotations provided by persons identifying as White as opposed to those identifying as Black or African American. The distance between their annotations measures .19 to .22 Likert scale points for offensiveness rating and .14 to .15 points for politeness rating. Additionally, the LLMs are progressively less accurate in reflecting the views of older individuals at offensiveness rating, with a .01 point increase in distance for each year of age. In terms of educational background, the models exhibit a greater discrepancy from those with less than a high school education compared to those with higher educational levels, ranging from .34 to .57 points for politeness rating. One possible explanation for these discrepancies by annotator sociodemographics is that the models’ training data may lack representation for these demographic groups.
@@
Notably, the analysis found no significant effects related to gender or occupational status. This could be because the LLMs do not consistently favor any particular group of people on these dimensions, or it could be because there are in fact no systematic differences in human annotation along these dimensions. Overall, while the statistically significant differences in annotation distances suggest some demographic biases, we emphasize that these effects are small compared to the full Likert scale range. In addition, the R2 Marginal values for these models (reported in Appendix A.1) are quite low, indicating that the sociodemographic categories explain little of the variability in human–LLM annotation differences. This complexity makes it challenging to establish a clear default persona for the LLMs. Therefore, we perform a second analysis in the following.
@@
Table 4: Sample sizes and mean distance scores of demographic-prompting predictions in comparison to no-info-prompting predictions for models GPT-4o and Claude at two rating tasks. Variables with very few cases (count $\leq$ 100) are shown in gray to indicate a lack or reliability of these numbers.
The Effect of Demographic Prompting. In a second approach, we examine how the inclusion of demographic information in the prompt (cf. second setting described in Section 3.5) influences the automatic annotation outputs of the models. We compare SD prompts with N prompts. This allows us to infer the demographics that the models defaults to.
Table 4 shows the results as differences in the prediction scores ($\Delta_{\mu}$). [^4] Note that there is an overlap in the age ranges (50–54, 54–59), a consequence of the survey question design in the original annotation task for the Popquorn dataset. We have preserved these original categories, despite this issue, in order to maintain the fine-grained distinctions they provide. For some demographic attributes, there are relatively few samples (“count” columns). This implies that there may not be sufficient statistical evidence to support the observed average distance values for certain categories. We focus our discussion on cases with $\geq$ 100 cases.
GPT-4o shows substantial differences in the scores for the offensiveness task across different gender attribute values: the predictions for the attribute value Non-Binary is substantially larger (.29) than for Male (.18) and Female (.20). This suggests that prompting the LLM to act as a non-binary individual has a more pronounced effect on its predictions, indicating that the male or female attributes are more aligned with its default persona. This observation is not consistent across tasks.
@@
We investigate whether the modifications to the model are more substantial for socio-demographic prompting compared to non-relevant additional information provided in the prompts. Specifically, we analyze how the models’ prediction scores for the two application tasks are affected when presented with placebo information (cf. third setting as given in Section 3.5).
We evaluate if these results are as pronounced as the observations based on Table 4 as described above. Overall, the results of the placebo prompting (P prompts) in comparison to N prompts do not reveal any notable differences for specific attribute values. The scores remain consistently stable across these comparisons. Therefore, we show the results only in the appendix, in Section A.3. Consequently, we conclude that the changes in model behavior are indeed more pronounced for socio-demographic prompting compared to the inclusion of non-relevant additional information.
@@
We investigate how task properties influence the role of demographic information in prompting and whether these patterns remain consistent across different application tasks. Our general conclusion indicates that there is no clearly distinguished default persona of the models for both tasks. Table 3 suggest that the results are in general consistent for the two tasks. However, we did also highlight some task-specific tendencies above at the analysis regarding RQ1. Notably, there are also differences in the absolute values of the prediction distances when comparing demographic prompting to no-information prompting across the two tasks. The average prediction difference associated with socio-demographic prompting of GPT-4o for offensiveness rating is calculated to be 0.19. In contrast, the average for politeness rating is substantially higher at 0.25. Similar results are evident for the model Claude (0.16 for offensiveness rating and 0.21 for politeness rating). This suggests that, in general, the LLMs are more influenced by demographic information at rating politeness than rating offensiveness.

### v3

@@ Abstract
Johannes Schäfer, Aidan Combs, Christopher Bagdon, Jiahui Li, Nadine Probol, Lynn Greschner, Sean Papay, Yarik Menchaca Resendiz, Aswathy Velutharambath, Amelie Wührl, Sabine Weber,  and Roman Klinger Fundamentals of Natural Language Processing, University of Bamberg, Germany roman.klinger@uni-bamberg.de
@@ 1 Introduction
Input Example:
If it was an American building it would collapse into itself before the fire spread to more than a few floors.
Annotation
[…]
Figure 1: Our objective is to identify which human demographic groups are mimicked by LLMs during subjective annotation tasks on text data.
@@
In some text [^1] We provide our code and model predictions on https://www.uni-bamberg.de/en/nlproc/resources/llms-default-demographics/. annotation tasks, it is feasible to obtain an aggregated ground truth label, for instance in named entity annotation (Yadav and Bethard, 2018) or semantic role labeling (Shi et al., 2020). In other tasks, perhaps even in the majority of tasks, it is more obvious that annotators’ traits influence the label assignments, for instance in sentiment annotation (Liu, 2012), emotion annotation (Klinger, 2023; Plaza-del Arco et al., 2024), or personality profiling (Neuman, 2015). The diversity of annotations, conditioned on annotators’ profiles, has been recognized as an important variable to consider, instead of aggregating all labels to an adjudicated score, which might not correspond to any of the annotators (Plepi et al., 2022).
@@
With the advent of (instruction-tuned) large language models (LLMs), automatic data annotation and zero or few shot predictions became more popular (Brown et al., 2020). However, language models do not provide the same diversity as human annotators do in a simple zero-shot setup, which can lead to a lower performance of the labeling process (Bagdon et al., 2024). To mitigate this problem, it is important to understand whether large language models exhibit a default persona when acting as annotators and how this can be controlled to ensure greater diversity in annotations. One concern is that large language models may reflect biases present in their training data, which can disproportionately emphasize certain viewpoints. This raises the issue of potential marginalization of minority perspectives in the outputs generated by these models, making it crucial to address this bias for equitable representation in annotations. One idea to address this limitation is socio-demographic prompting (Muscato et al., 2024), where we guide the model by specifying characteristics in the prompt, like age, gender, ethnicity or socio-economic status.
@@
Previous research has explored the effects of using prompts informed by demographic or cultural contexts on model predictions, but did not find consistent patterns (Beck et al., 2024; Mukherjee et al., 2024). We build on top of this work and particularly contribute in two directions of socio-demographic prompting. We study if large language models default to a particular demographic, i.e., we evaluate if the LLM-based predictions more closely align with those of people of a particular demographic when not being conditioned. Our general objective is depicted in Figure 1. Additionally, we utilize placebos as suggested by Mukherjee et al. (2024) – irrelevant information that should not affect the model’s output – to compare the impact of demographic prompts and evaluate the stability of the model’s predictions. More concretely, we answer the following questions.
@@
The remainder of this paper is structured as follows: We review related work in Section 2 and explain our data set and task choices as well as prompt setup in Section 3. In Section 4, we present the results of our experiments and conclude including a discussion of possible future work in Section 5.
@@ 2 Related Work
This paper relates to various areas which we review in the following, namely biases in large language models, perspectivism, and (socio-demographic) prompting of LLMs.
@@ 2.1 Inherent Knowledge and Biases in LLMs
Large language models may be understood as databases that store information encoded in the training data (Petroni et al., 2019). These data are stored in a probabilistic manner, which allows models to show sometimes unexpected generalization capabilities beyond the original instructions available in the training data, i.e., emergent abilities (Wei et al., 2022). It has also been shown that prompting models can adapt parameters during inference, similar to how backpropagation works (von Oswald et al., 2023). Therefore, it is reasonable to view prompting as a form of programming, as it involves manipulating model behavior by crafting specific inputs to achieve desired outputs (Beurer-Kellner et al., 2023).
@@ 2.2 Perspectivism
Natural language processing for a long time focused on seemingly objective tasks such as parsing or named entity recognition. Therefore, it has been a standard approach to adjudicate annotations into a single gold standard (Stubbs, 2011). To do so, a set of methods for aggregation that considers the distributions and disagreements have been developed (Paun et al., 2018). More recently, research moved towards acknowledging the importance of disagreements. Plank (2022) has been one of the first to discuss challenges and potential approaches to this problem. Recent work looked into methods to consider disagreements (Fleisig et al., 2024) and discusses ethical considerations (Valette, 2024). Nowadays, this field perspectivism developed dedicated shared tasks (Uma et al., 2021) and workshops (Abercrombie et al., 2024).
Disagreement is related to confidence of annotators (and models), an aspect that has received some attention. Baumler et al. (2023) made use of this property in the context of active learning, to understand which instances require annotations by multiple people in order to understand the disagreements. Troiano et al. (2021) and Baan et al. (2024) showed that annotators’ own confidence predicts inter-annotator agreement scores.
To understand disagreement in labeling better, more and more corpora are being published with more detailed information about the annotators. Troiano et al. (2023) annotated event reports for emotions and appraisals and collected demographic information, personality, and current emotional state of both the person who lived through the event and multiple annotators that read the event description. Plepi et al. (2022) studied the role of demographics, automatically extracted from the data, on the perception of social norms. Bizzoni et al. (2022) discuss the role of individual differences on the judgement of literary quality. Romberg (2022) integrates perspectivism in argument mining, by making explicit the subjective nature of argument interpretation. May et al. (2024) study the effect of demographics on the role of numbers in social judgements. Frenda et al. (2023) study how the perception of irony varies by nationality, employment status, student status, ethnicity, age, and gender. Sachdeva et al. (2022) measure different aspects of hate speech which include sentiment, disrespect, insult, attacking/defending, humiliation, inferior/superior status, dehumanization, violence, genocide, and a 3-valued hate speech benchmark label. They study these variables under the condition of identity target groups and annotator demographics. Xu et al. (2023) look into disagreement of legal case outcome differences. Next to subjective tasks, perspectivism has also been considered in seemingly objective tasks, for instance in named entity recognition (Peng et al., 2024) and natural language inference (Gruber et al., 2024).
@@ 2.4 Socio-Demographic Prompting
Automatic annotation with language models is not a replacement for human annotation. Humans have previous world knowledge, experiences, and perspectives on a matter that differ individually. LLMs have difficulties replicating these differences, causing issues in annotation. For instance, Bagdon et al. (2024) show that the diversity of annotations that is beneficial in comparative annotations by multiple people cannot be straightforwardly replaced by multiple runs of a language model. Lee et al. (2023) focus on this aspect in particular and find poor alignment of the distribution of labels predicted by LLMs with human annotations on natural language inference tasks.
@@
Additionally, Sun et al. (2025) highlight that most LLMs show demographic biases in subjective judgment tasks, favoring perceptions from White participants over those from Asian or Black participants. Hu and Collier (2024) find that incorporating persona variables in LLM prompting improves model predictions slightly, especially in conditions of significant annotator disagreement. Movva et al. (2024) reveal that while GPT-4 shows reasonable alignment with human assessments of safety, there are significant demographic disparities in how well it correlates with different annotator groups.
@@ 3.1 Data Sets
We chose the Popquorn data for our experiments (Pei and Jurgens, 2023). In contrast to data used by other previous research we are aware of, these data have been particularly sampled for the study of annotators’ properties and their impact on annotation tasks. Therefore, these data render themselves as a straightforward choice also for an LLM-based analysis. The original data stem from the Ruddit corpus (Hada et al., 2021) which was originally annotated via best–worst scaling. This may lead to different label frequencies than rating scale annotations and is considered as not appropriate for skewed distributions (Finn and Louviere, 1992; Louviere et al., 2015). The authors of Popquorn reannotated the data with 1–5 rating scales, which we adopt for our annotation setup. Further, it is noteworthy that the creators of the data set removed annotators with particularly low annotation competency according to Mace (Hovy et al., 2013), to filter out potential random noise.
The Popquorn data set consists of 45,000 annotations from 1,484 annotators with information on their gender, race, age, occupation and education. We use data from the subjective partitions for the offensiveness and politeness rating tasks. To equalize the number of annotations per instance, we randomly sample three annotations for each instance. [^2] The decision to retain three annotators per instance was a deliberate choice to standardize our methodology, as the original dataset has varying numbers of annotators. We aimed to ensure consistency across our analysis and view the balance we achieve as a necessity for the comparison to non-demographic prompting, which always produces one prediction per instance. With our method, we avoid a skewed distribution towards instances with a larger number of annotators in demographic prompting. We have checked whether the distribution is skewed due to our sampling: For the offensiveness data, our sample consists of 35% of the overall annotations from Popquorn. As for the distribution of each demographic attribute in our sample: It comprises 31%-38% of the annotations per attribute (standard deviation: 1%). For the politeness data, our sample consists of 45% of the overall annotations. As for the distribution of each demographic attribute in our sample: It comprises 34%-56% of the annotations per attribute (standard deviation: 4%; the wide range is due to the fact that some attributes being extremely rare in the original dataset, leading to outliers). Our sample maintains a certain degree of representativeness of the demographics in Popquorn. When sampling, we omit the annotations where the annotator preferred not to disclose demographic information. We exclude data instances with fewer than three annotations remaining. The two subsets of the sample which we use in our experiments consist of 4,500 annotations of 1,500 instances for offensiveness ratings and 11,151 annotations of 3,717 instances for politeness ratings.
@@
Table 2: Sets of values for placebo attributes used in P prompts in our experiments.
3.2 Prompt Setup
We use three different prompt setups, namely socio-demographically informed (SD) prompts, placebo-informed (P) prompts, and prompts without additional information (N prompts). We show the templates used for these prompts in Table 1. The input to the LLM consist of a concatenation of the system message with the respective text to be classified. The values for the demographic attributes are taken from the demographic data of the annotators included in the sample. The sets of values for the placebo attributes are displayed in Table 2. During prompting, each placebo attribute value is randomly sampled from the respective set.
@@ 3.3 Model Choice and Access
We use two LLMs, namely GPT-4o (OpenAI, 2024b) and Claude (Anthropic, 2024). We access both models via their respective APIs with default hyperparameters. The total cost for using GPT-4o was \$54 and the cost for using Claude was \$60.
@@ 3.4 LLM Output Parsing
We transform the texts generated by the LLMs for the annotation tasks into a defined label set of five categories, corresponding to the 1–5 rating scale for each of the two tasks. We use the Langchain StrOutputParser [^3] https://api.python.langchain.com/en/latest/output_parsers/langchain_core.output_parsers.string.StrOutputParser.html to interpret the output from GPT-4o as one of the designated labels, while we use the output from Claude directly.
This process does not successfully parse the output in all cases. We encountered 22 error cases with GPT-4o and three with Claude. Of the failed cases with GPT-4o, 14 were due to the input text from the original data set being in Polish instead of English. This resulted in the model predicting labels in Polish, despite the prompt specifying an English-only label set. The remaining eight cases involved instances with very short text or questions, which the model misinterpreted. In these cases, GPT-4o asked for further input instead of performing the intended classification task.
For Claude, the three instances of failure involved the names of famous actors in the text. This triggered the model’s copyright protection protocols, preventing it from reproducing or paraphrasing potentially copyrighted content.
Given the total number of instances analyzed in our experiments, this failure rate is insubstantial. Consequently, we disregard the output in the cases where parsing was unsuccessful and assign the labels “not offensive” or “not polite”. [^4] Our decision to label such instances with the negative class, rather than removing them, was made to maintain comparability across different experimental settings. Removing instances due to a single failed label would distort the consistency of the number of variants of prompts per text instance.
@@ 3.5 Evaluation Settings
We conduct various analyses of the labels generated by the LLMs, utilizing the different prompt setups described in Section 3.2. First, we examine the extent to which LLM predictions align with the judgments of different human annotators by comparing the outputs generated from non-demographic (N) prompts to the human annotations available for each instance. Second, we examine the impact of socio-demographic prompting by comparing the models’ automatic annotations produced with SD prompts to those produced with N prompts. In this analysis, we assess the differences in the LLMs’ annotations when demographic data is included versus when it is omitted. Thirdly, we investigate the effect of placebo prompting in a similar manner by comparing the automatic annotations generated with P prompts to those created with N prompts.
@@ 4 Results
In this section, we present the results of our experiments based on the annotations generated by the LLMs for two labeling tasks. Each of our research questions is addressed with particular results derived from the various experiments conducted. Detailed results for individual prompts show that the two N prompts behave similarly (see Section A.2). We also observed this for the N and SD prompt templates. Thus, we here report all results as an average over the two respective prompt templates for the different prompt types.
@@
We approach the identification of the default persona of the models in two different ways following the first two settings as described in Section 3.5. The following sections present and discuss the results of these analyses.
@@
Which socio-demographic attributes of human annotators does an LLM inherently mimic in the absence of explicit information (N prompts)? Table 3 shows the results for this analysis which corresponds to the first setting described in Section 3.5. We combine socio-demographic attributes that are represented in only a few cases. The table specifies the reference categories (ref.) used for each of the categorical variables.555 The reference categories for the categorical variables occupation and education are chosen according to the recommendations of Johfre and Freese (2021). For race and gender, their scheme does not result in unambiguous recommendations, so we choose the categories we expect to lie closest to the models’ defaults: male for gender and white for race.
@@
Table 3 shows the coefficients of regression models predicting absolute values of the distance between LLM annotations (N prompts) and annotations provided by human annotators for the same instance. Independent variables are the socio-demographic characteristics of the human annotators. We report results separately for the two models (GPT-4o and Claude) and two classification tasks (offensiveness and politeness rating). The reported coefficients can be interpreted as the effect of particular human demographic categories on the distance between human and LLM annotations (change in distance between the category in question and the reference category). Positive coefficients indicate that the respective LLM is further from the human annotators in that demographic category than human annotators in the reference category, i.e., less accurate. From this we can follow that the categories with positive, statistically significant coefficients are those categories which the respective model does not default to. In contrast, negative coefficients indicate that the LLM is closer to the human annotators in the given demographic category than it is to the human annotators in the reference category; i.e., more accurate. Categories with negative and statistically significant coefficients are those that are nearer to the model’s default than the reference category.
@@
The analysis of coefficients (see Table 3) reveals significant biases in the LLMs’ predictions based on demographic factors, particularly concerning race and age. Specifically, the models demonstrate a tendency to align more closely with annotations provided by persons identifying as White as opposed to those identifying as Black or African American. The distance between their annotations measures .19 to .22 Likert scale points for offensiveness rating and .14 to .15 points for politeness rating. Additionally, the LLMs are progressively less accurate in reflecting the views of older individuals at offensiveness rating, with a .01 point increase in distance for each year of age. In terms of educational background, the models exhibit a greater discrepancy from those with less than a high school education compared to those with higher educational levels, ranging from .34 to .57 points for politeness rating. One possible explanation for these discrepancies by annotator sociodemographics is that the models’ training data may lack representation for these demographic groups.
@@
Notably, this analysis does not show any significant effects related to gender or occupational status. This could be because the LLMs do not consistently favor any particular group of people on these dimensions, or it could be because there are in fact no systematic differences in human annotation along these dimensions. Overall, while the statistically significant differences in annotation distances suggest some demographic biases, we emphasize that these effects are small compared to the full Likert scale range. In addition, the R2 Marginal values for these models (reported in Appendix A.1) are quite low, indicating that the sociodemographic categories explain little of the variability in human–LLM annotation differences. This complexity makes it challenging to establish a clear default persona for the LLMs. Therefore, we perform a second analysis in the following.
@@
Table 4: Sample sizes (count) and mean distance scores of demographic-prompting (SD prompts) predictions in comparison to predictions with N prompts for models GPT-4o and Claude at two rating tasks. Variables with very few cases (count $\leq$ 100) are shown in gray to indicate a lack or reliability of these numbers.
The Effect of Demographic Prompting. In a second approach, we examine how the inclusion of demographic information in the prompt (cf. second setting described in Section 3.5) influences the automatic annotation outputs of the models. We compare the outputs generated with SD prompts to those with N prompts. We assess the differences in the LLMs’ annotations with and without the inclusion of demographic data, which enables us to infer the demographics closest to the models’ default.
Table 4 shows the results as differences in the prediction scores ($\Delta_{\mu}$). [^6] Note that there is an overlap in the age ranges (50–54, 54–59), a consequence of the survey question design in the original annotation task for the Popquorn dataset. We have preserved these original categories, despite this issue, in order to maintain the fine-grained distinctions they provide. For some demographic attributes, there are relatively few samples (“count” columns). This implies that there may not be sufficient statistical evidence to support the observed average distance values for certain categories. We focus our discussion on cases with count $\geq$ 100.
GPT-4o shows substantial differences in the scores for the offensiveness task across different gender attribute values: The prediction differences for Non-Binary (.29) are substantially larger than those for Male (.18) and Female (.20). This suggests that prompting the LLM to act as a non-binary individual has a more pronounced effect on its predictions, indicating that the male or female attributes are more aligned with its default persona. This observation is not consistent across tasks.
@@
The discrepancy between the results from the analysis displayed in Table 3 to those from the analysis displayed in Table 4 are not a contradiction. These analyses are based on distinct interpretations of the default persona of the models. Table 3 indicates a poor representation of some demographics in the models’ responses. Conversely, Table 3 highlights that, when prompted with specific demographics, the output was marginally closer to the default, suggesting that the models’ interpretation of these demographic attributes influences its behavior only slightly. Thus, the model exhibits a measurable lack of alignment with certain demographics while it simultaneously demonstrates minimal variation in its behavior when prompted to act as a person with those demographics.
@@
We investigate whether the modifications to the model are more substantial for socio-demographic prompting compared to non-relevant additional information provided in the prompts. Specifically, we analyze how the models’ predictions for the studied tasks are affected when presented with placebo information (cf. third setting from Section 3.5).
We evaluate if these results are as pronounced as the observations based on Table 4 as described above. Overall, the results of the placebo prompting (P prompts) in comparison to N prompts do not reveal any notable differences for specific attribute values (see Section A.3). The scores remain consistently stable across these comparisons. Consequently, we conclude that the changes in model behavior are indeed more systematic for prompting with specific socio-demographic attributes than when irrelevant additional information is included.
The changes resulting from placebo prompting in comparison to N prompts appear to be substantial, with some discrepancies even surpassing those related to socio-demographic prompting. This pattern is more pronounced within the offensiveness task. The discrepancies arise from the analysis that focuses on the absolute values of the differences, regardless of their direction, thus capturing notable fluctuations that may not reflect a consistent trend.
@@
We investigate how task properties influence the role of demographic information in prompting and whether patterns remain consistent across the tasks. Our general conclusion indicates that there is no clearly distinguished default persona of the models for both tasks. Table 3 shows that results are in general consistent across the two tasks. However, the analysis regarding RQ1 also highlights some task-specific tendencies. Notably, there are also differences in the prediction distances when comparing demographic prompting to no-information prompting across the two tasks. The average prediction difference associated with socio-demographic prompting of GPT-4o for offensiveness rating is calculated to be 0.19. In contrast, the average for politeness rating is substantially higher at 0.25. Similar results are evident for the model Claude (0.16 for offensiveness rating and 0.21 for politeness rating). This suggests that, in general, these LLMs are more influenced by demographic information at rating politeness than rating offensiveness.